OpenStudy (anonymous):

Find the amount of decrease if 68 is decreased by 15%

OpenStudy (anonymous):

Ok so first you have to find what 15% of 68 is 15%=15/100 you are trying to find what x/68 is cross multiply 15(68)=1020 100(x)=100x 1020/100 x=10.2 so now you know that 15% of 68=10.2 simply subtract 10.2 from 68 to find your answer Your answer should be: 57.8

OpenStudy (anonymous):

actually the amount of decrease is 10.2

OpenStudy (anonymous):

Awesome thank you.

OpenStudy (anonymous):

No problem! Anytime:)

OpenStudy (anonymous):

Might just have to hire you on lol ill pay lol

OpenStudy (anonymous):

Ha ha no but I am willing to help anytime